Suppose an economy has two years worth of data, years 1 and 2. Suppose there are also two goods, bread and corn. Suppose in year one fifty units of corn are sold at a price of 1 $ and 20 units of bread are sold at a price of 2 $. Suppose in year two, 60 units of corn are sold at a price of 1.5 $ and 80 units of bread are sold at a price of 2.05 $. For all three measures, compute the GDP deflator. What are the implications for these measures for the amount of inflation in this economy. [Note, you should have five candidate measures]. What about the amount of economic growth? Suppose that corn in year 2 is twice as valuable as corn in year 1 to consumers in terms of their enjoyment from its consumption (or in terms of its nutritional value). How would your above calculations change numerically if we want real GDP growth to include improvements in quality?
Answer #1

Solution

From the information given in question

Year Q​​​​​​C P​​​​​​C Q​​​​​​B P​​B
1 50 1 20 2
2 60 1.5 80 2.05

​​Calculation

Nominal GDP (Year1) = 50*1+20*2=90

Nominal GDP (Year2) = 60*1.5+80*2.05=90+164=254

Taking Year1 prices as base year

Real GDP (1) = 50*1+20*2=90

Real GDP (2) = 60*1+80*2=220

GDP Deflator (Year 1)

=Nominal GDP(Year 1)/Real GDP (Year)*100

=90/90*100= 100

GDP Deflator (Year 2) =254/220*100=115.45

Economic Growth in %

= Real GDP(Year2)- Real GDP(Year1) /Real GDP (Real1)*100

= 220-90/90*100= 130/90*100 =144.44

If corn in year 2 is twice valuable than year 1 then there won't be any change in the above calculation until we change the base year